OpenStudy (anonymous):

I have an answer can anyoe tell me what they got. Find the absolute maximum and minimum for y=x^2+(2/x) on 0.3 < or equal to x < or equal to 4

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

for max I got-4 min I got 8

OpenStudy (anonymous):

Is that correct

OpenStudy (anonymous):

now, check the domain :)

OpenStudy (anonymous):

\[0.3\le x\le4\]

OpenStudy (anonymous):

sorry but I a a little confused

OpenStudy (anonymous):

hold it.. thats not it!!!

OpenStudy (anonymous):

\[y=x^2+2x^{-1}\\ y'=2x-2x^{-2}=0\\\implies2x-{2\over x^2}=0\implies x^3-1=0\implies\boxed{x=1} \] there is only one critical value.. (note that x=0 is not in the domain of the function)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the the max is -1

OpenStudy (anonymous):

how did you get the "-1"??

OpenStudy (anonymous):

sorry I meant1

OpenStudy (anonymous):

to check if it is a maxima or a minima, we need the second derivative.. so, find the second derivative

OpenStudy (anonymous):

ok it would be 2x+4x^-1=0

OpenStudy (anonymous):

\[y'=2x-2x^{-2}\\y''=?\]

OpenStudy (anonymous):

it would be \[-\frac{ 12 }{ x^4}\]

OpenStudy (anonymous):

 \[y''=2+4x^{-3}\]